OpenStudy (anonymous):

On a map of New Mexico, 1 cm represents 45 km. The distance from Albuquerque to Santa Fe is 2 cm. Find the actual distance. (1 point) 90 km 22.5 km 15 km 60 km

OpenStudy (phoenixfire):

1cm=45km The distance is 2cm. 2cm=2*1cm=2*45km=?

OpenStudy (anonymous):

Thanks :D